Let S= {(x,y)|x,y € R}. How do we show that S is a ring with identity with the operations defined by (x,y) +(u,v) = (x+u, y+v) and (x,y)(u,v) = (xu-yv,xv+yu)?


1
Expert's answer
2021-04-14T02:39:04-0400

(0,0) "\in" S


Since 0 "\isin" R


Additive identity is (0,0)

Considering (x,y) "\isin" S


(x,y)+(0,0)=(x+0,y+0)=(x,y)


"\therefore" Additive identity exists

If e=(r,s) is the multiplicative identity then,

(x,y)(r,s)=(x,y)

=(xr-ys,xs+yr)


Solving this we get r=1 and s=0


"\therefore" the multiplicative identity s (1,0)



hence the identity element exists